y= x^3 - x^2 + x + 5, find  dy/dx.


y=4x^3 + 7x^2 - 4x + 8, find  dy/dx.


s=17t^3 - 5t^2 - 5t - 3, find  ds/dt.


1
Expert's answer
2021-04-15T07:19:49-0400

"dy/dx = \frac d{dx}(x^3 - x^2 + x + 5) = 3x^2 -2x + 1"


"dy/dx = \frac d{dx}(4x^3 + 7x^2 - 4x + 8) = 12x^2 + 14x - 4"


"ds/dt = \frac d{dt}(17t^3 - 5t^2 - 5t - 3) = 51t^2 - 10t - 5"